OpenStudy (anonymous):

What is the midpoint of the line segment with endpoints (5, -2) and (1, -6)?

OpenStudy (anonymous):

\[m-point =(\frac{5+1}{2}, \frac{-2-6}{2}) =(\frac{6}{2}, \frac{-8}{2})=(3,-4)\]

OpenStudy (anonymous):

tanks

OpenStudy (anonymous):

when ur finding midpoint of end points all you gotta do is add x and y and divide them by 2